# HGCLIP: Exploring Vision-Language Models with Graph Representations for Hierarchical Understanding

### v1

@@ Appendix B Method Introduction
Calculating Logits for Hierarchical Classification: The calculation of the predictive probabilities for each hierarchical level primarily involves three methods [6]. The first method, as depicted in Figure 9, describes a classifier that functions on a per-level basis, designated as $h$ individual $\{K_{i}\}_{i=1}^{h}$-way classifiers. The model utilizes information about the label hierarchy by explicitly predicting a single label per level for a given image. However, this computational approach cannot be applied to methods based on CLIP. This is because VLMs determine the predicted class by calculating the similarity between the features of the image and those of each text prompt, and the model does not possess multiple linear layers to correspond to the prediction results of each hierarchical level separately. Second, Figure 10 illustrates a model configured as a hierarchy-agnostic classifier. This particular multi-label classifier operates without incorporating any knowledge of an explicit hierarchical structure within the labels. Our experiments are based on this computational method to calculate the predictive probabilities at each hierarchical level. Third, Figure 11 presents a model schematic for the marginalization approach, wherein, rather than assigning a discrete label to each level, the model generates a probabilistic distribution across the terminal nodes of the hierarchical structure. The probability associated with non-terminal nodes is computed by marginalizing across their immediate offspring. This marginalization method not only captures the connections of the various nodes within the hierarchy but also acknowledges the presence of $h$ levels within the label hierarchy. However, employing this method to compute the predictive probabilities of superclasses yields poor results. Detailed comparative results are presented in Table 8.
Figure 9: Model schematic for the per-level classifier.
@@
Figure 10: Model schematic for the hierarchy-agnostic classifier.
Figure 11: Model schematic for the marginalization method.
@@
Appendix C Ablative Analysis
Hyperparameter Settings: In Figure 12, We conduct ablation on $\lambda_{1}$ and $\lambda_{2}$ hyper-parameters of logits with structured features for CIFAR-100 dataset. Overall, an increase in $\lambda_{1}$ can enhance performance across multiple hierarchical levels, indicating that the textual structured information is rich and effective. On the other hand, with variations in $\lambda_{2}$, the performance across various levels peaks around $\lambda_{2}$ = 0.2. This reflects that structured image features provide more effective information, which in turn improves model performance.
Analysis of Noisy Hierarchies Caused by LLMs: Taking Food-101 as an example, we conduct two types of comparative experiments utilizing hierarchical labels obtained from the LLMs (i.e., ChatGPT [33]). The first experiment involves repeating the same input and testing the various hierarchical labels generated by the LLMs. Despite LLMs generating suboptimal hierarchical labels, in Table 10, our method still enhances model performance based on these noisy hierarchical labels. This demonstrates the strong robustness and generalizability of HGCLIP. The second experiment involves generating hierarchical labels of different depths and conducting experiments based on these hierarchies. In Table 10, across varying numbers of levels in label hierarchies, HGCLIP improves model performance, even though some categories in the label hierarchies may be overly broad or detailed. Through the validation of these two types of experiments, our approach effectively utilizes the label hierarchy structure, thus improving hierarchical classification performance.
@@ Appendix C Ablative Analysis
Table 9: Performance based on hierarchical labels obtained from LLMs given the same input.
Method Acc. %
$l_{1}$ $l_{2}$
CLIP 61.38 85.53
HGCLIP 91.12 88.73
[…]
HGCLIP 90.67 87.20
Method Acc. %
@@
CLIP 61.38 85.53 -
@@
Table 9: Performance based on hierarchical labels obtained from LLMs given the same input.
Table 10: Comparative results based on hierarchical labels of different numbers of levels queried by LLMs.
Appendix D Results Presentation
Detailed results on subpopulation shift benchmark: In Table 12, we present the detailed results for the subpopulation shifts benchmark illustrated in Figure 5.
t-SNE visualization: In Figure 3, we present the visualization of image embeddings, which demonstrates that our method effectively separates categories across multiple granularities. However, it is hard to observe whether the features of certain categories with hierarchical relationships are separable. To intuitively understand the role played by the graph representations leveraging class hierarchy, we visualize the features of one single coarse category and its corresponding five fine-grained categories randomly selected from CIFAR-100 in Figure 13. It is evident that the features of the fine-grained categories from HGCLIP are more distinct compared to both CLIP and MaPLe [22]. This can be attributed to the utilization of graph representations that integrate hierarchical relationships between categories. Similarly, the features at the coarse-grained level also benefit. CLIP does not focus on the intra-class arrangement of samples within the feature space, as long as these samples are proximate to each other and are separable from samples of different categories. This can lead to a chaotic distribution of subcategories within the same category region. In contrast, HGCLIP not only retains the attribute features of the coarse categories but also reduces the intra-class variance, achieving enhanced generalization.
Analysis with CHiLS [32] and Hierarchy-CLIP [11]: These two works are among the most recent and earliest to combine label hierarchies with VLMs. Specifically, CHiLS utilizes the subcategories of each label to enrich the class name itself, which is effective in cases where datasets may contain rich underlying structures but have information-poor class labels. However, this approach experiences a notable performance decline when applied to some fine-grained datasets. This is because the class labels in fine-grained datasets are already very specific, and forcibly adding more specialized and specific subcategories can overwhelm the model, leading to an inability to match images with the correct text. Building upon this, Hierarchy-CLIP enriches class labels to a greater extent by adding both the superclass and subclasses to the class name after initially identifying the top-5 class candidates. This effectively compensates for the shortcomings of CHiLS. Nevertheless, this method is still training-free and cannot be maximally adapted to downstream datasets. Fine-tuning methods based on CLIP, whether prompt tuning [54, 55, 22, 20] or feature adaptation [10, 53, 14], do not integrate label hierarchies. Therefore, the introduction of HGCLIP effectively addresses these limitations. It is trainable and can be easily transferred to datasets of different semantic granularities (regardless of the availability of label hierarchies), and it also enhances the classification accuracy of each semantic hierarchical category in the dataset. In Table 11, we compare the performance of our method with that of these two approaches. Regarding the CHiLS experiment, as their paper indicates that it is tested only on the 20 classes of CIFAR-20 and 101 classes of Food-101, to maximize the representation of its effect, we follow its setup and do not test it on the remaining hierarchical categories. The results show an improvement on the generic image dataset, but a decline on the fine-grained food dataset, consistent with our previous analysis. In the Hierarchy-CLIP experiment, it also exhibits commendable performance, showing improvements across various hierarchical levels. This is largely attributed to its use of both superclasses and subclasses of labels for expansion. Our method not only enhances classification performance at each level but also can be conveniently transferred to downstream datasets.
@@ Appendix D Results Presentation
CLIP [36] ICML’21 43.22 66.57 61.38 85.53
CHiLS [32] ICML’23 52.58 - - 83.18
Hierarchy-CLIP [11] CVPR’23 58.61 67.39 63.74 86.98
@@
Method Entity13 Entity30 Living17 NonLiving26
Base Novel Base Novel Base Novel Base Novel
CLIP 68.55 68.62 70.32 67.37 95.00 88.24 73.04 76.42
CoOp 94.22 85.99 89.83 78.35 95.59 88.77 85.23 76.69
CoCoOp 95.06 85.23 91.52 78.65 96.35 87.77 91.65 77.15
MaPLe 97.03 84.95 93.75 78.63 97.47 87.24 93.62 74.85
PromptSRC 93.95 85.59 89.30 79.35 95.53 89.00 86.85 79.77
HGCLIP 97.91 86.10 93.99 79.67 98.12 88.93 94.62 79.86
Table 12: Detailed performance (%) of different methods on the four BREEDS datasets [39] to measure robustness to subpopulation shift.

### v3

@@ D.1 Calculating Logits for Hierarchical Classification
The calculation of the predictive probabilities for each hierarchical level primarily involves three methods Dhall et al. (2020). The first method, as depicted in Figure 8(a), describes a classifier that functions on a per-level basis, designated as $h$ individual $\{K_{i}\}_{i=1}^{h}$-way classifiers. The model utilizes information about the label hierarchy by explicitly predicting a single label per level for a given image. However, this computational approach cannot be applied to methods based on CLIP. This is because VLMs determine the predicted class by calculating the similarity between the features of the image and those of each text prompt, and the model does not possess multiple linear layers to correspond to the prediction results of each hierarchical level separately. Second, Figure 8(b) illustrates a model configured as a hierarchy-agnostic classifier. This particular multi-label classifier operates without incorporating any knowledge of an explicit hierarchical structure within the labels. Our experiments are based on this computational method to calculate the predictive probabilities at each hierarchical level. Third, Figure 8(c) presents a model schematic for the marginalization approach, wherein, rather than assigning a discrete label to each level, the model generates a probabilistic distribution across the terminal nodes of the hierarchical structure. The probability associated with non-terminal nodes is computed by marginalizing across their immediate offspring. This marginalization method not only captures the connections of the various nodes within the hierarchy but also acknowledges the presence of $h$ levels within the label hierarchy. However, employing this method to compute the predictive probabilities of superclasses yields poor results. Detailed comparative results are presented in Table 8.
@@
Figure 8: (a) Model schematic for the per-level classifier. (b) Model schematic for the hierarchy-agnostic classifier. (c) Model schematic for the marginalization method.
Figure 9: Illustrations of GPT-generated hierarchical labels.
D.2 GPT-mapped Hierarchical Labels
This is an illustrative diagram of hierarchical labels generated by ChatGPT OpenAI (2023). As demostrated in Figure 9, the prompts include all original class names and the desired number of hierarchy levels. ChatGPT will output the specified hierarchical labels.
@@ D.2 GPT-mapped Hierarchical Labels
Appendix E Ablative Analysis
@@
E.1 Hyperparameter Settings
In Figure 10, We conduct ablation on $\lambda_{1}$ and $\lambda_{2}$ hyper-parameters of logits with structured features for CIFAR-100 dataset. Overall, an increase in $\lambda_{1}$ can enhance performance across multiple hierarchical levels, indicating that the textual structured information is rich and effective. On the other hand, with variations in $\lambda_{2}$, the performance across various levels peaks around $\lambda_{2}$ = 0.2. This reflects that structured image features provide more effective information, which in turn improves model performance.
E.2 Analysis of Noisy Hierarchies Caused by LLMs
Taking Food-101 as an example, we conduct two types of comparative experiments utilizing hierarchical labels obtained from the LLMs (i.e., ChatGPT OpenAI (2023)) (in Table 9). (i) The first experiment involves repeating the same input and testing the various hierarchical labels generated by the LLMs. Despite LLMs generating suboptimal hierarchical labels, our method still enhances model performance based on these noisy hierarchical labels. This demonstrates the strong robustness and generalizability of HGCLIP. (ii) The second experiment involves generating hierarchical labels of different depths and conducting experiments based on these hierarchies. Across varying numbers of levels in label hierarchies, HGCLIP improves model performance, even though some categories in the label hierarchies may be overly broad or detailed. These experiments involve the performance disparity between GPT-mapped hierarchical labels and the officially provided annotation hierarchy, including inconsistencies in the labels or hierarchical levels. Through the validation of these experiments, our approach effectively utilizes the label hierarchy structure, thus improving hierarchical classification performance.
Type Method Acc. %
@@ E.2 Analysis of Noisy Hierarchies Caused by LLMs
Ground-Truth Hierarchical Labels CLIP 61.38 85.53 -
MaPLe 88.16 88.04 -
@@
Table 9: Results based on GPT-mapped and annotated hierarchical labels.
Appendix F Results Presentation
Our comparison methods are primarily divided into visual-only fine-grained visual classification methods and CLIP-based fine-tuning methods. The former is a uni-modal approach, utilizing only visual features and implementing hierarchical classification through multiple classification layers. We implement state-of-the-art methods from the past three years for comparison Du et al. (2020); Chang et al. (2021); Zhao et al. (2021); Liu et al. (2022); Xu et al. (2023). The latter comprises CLIP-based multi-modal methods, categorized into three types: zero-shot, feature adaptation, and prompt tuning. For zero-shot, we consider the only two methods Ge et al. (2023); Novack et al. (2023) combining class hierarchy and CLIP available to date in Appendix F.2. We implement several popular methods of feature adaptation Gao et al. (2023); Zhang et al. (2022); Guo et al. (2023); Song et al. (2023). Prompt tuning is the most common method for fine-tuning CLIP. We compare several popular methods from the past three years Zhou et al. (2022b, a); Jia et al. (2022); Yao et al. (2023); Khattak et al. (2023a, b).
Furthermore, it is important to note for experimental fairness that we are the first to migrate CLIP to hierarchical image classification, so there are no directly comparable methods that utilize both class hierarchy and CLIP. Currently, there are only two zero-shot methods that simultaneously utilize CLIP and class hierarchy. They are training-free, but we also compare them.
@@ Appendix F Results Presentation
F.1 t-SNE visualization
In Figure 3, we present the visualization of image embeddings, which demonstrates that our method effectively separates categories across multiple granularities. However, it is hard to observe whether the features of certain categories with hierarchical relationships are separable. To intuitively understand the role played by the graph representations leveraging class hierarchy, we visualize the features of one single coarse category and its corresponding five fine-grained categories randomly selected from CIFAR-100 in Figure 11. It is evident that the features of the fine-grained categories from HGCLIP are more distinct compared to both CLIP and MaPLe Khattak et al. (2023a). This can be attributed to the utilization of graph representations that integrate hierarchical relationships between categories. Similarly, the features at the coarse-grained level also benefit. CLIP does not focus on the intra-class arrangement of samples within the feature space, as long as these samples are proximate to each other and are separable from samples of different categories. This can lead to a chaotic distribution of subcategories within the same category region. In contrast, HGCLIP not only retains the attribute features of the coarse categories but also reduces the intra-class variance, achieving enhanced generalization.
@@
CLIP Radford et al. (2021) 43.22 66.57 61.38 85.53
CHiLS Novack et al. (2023) 52.58 - - 83.18
Hierarchy-CLIP Ge et al. (2023) 58.61 67.39 63.74 86.98
@@
F.2 Analysis with CHiLS Novack et al. (2023) and Hierarchy-CLIP Ge et al. (2023)
These two works are among the most recent and earliest to combine label hierarchies with VLMs. Specifically, CHiLS utilizes the subcategories of each label to enrich the class name itself, which is effective in cases where datasets may contain rich underlying structures but have information-poor class labels. However, this approach experiences a notable performance decline when applied to some fine-grained datasets. This is because the class labels in fine-grained datasets are already very specific, and forcibly adding more specialized and specific subcategories can overwhelm the model, leading to an inability to match images with the correct text. Building upon this, Hierarchy-CLIP enriches class labels to a greater extent by adding both the superclass and subclasses to the class name after initially identifying the top-5 class candidates. This effectively compensates for the shortcomings of CHiLS. Nevertheless, this method is still training-free and cannot be maximally adapted to downstream datasets. Fine-tuning methods based on CLIP, whether prompt tuning Zhou et al. (2022a, b); Khattak et al. (2023a); Jia et al. (2022) or feature adaptation Gao et al. (2023); Zhang et al. (2022); Guo et al. (2023), do not integrate label hierarchies. Therefore, the introduction of HGCLIP effectively addresses these limitations. It is trainable and can be easily transferred to datasets of different semantic granularities (regardless of the availability of label hierarchies), and it also enhances the classification accuracy of each semantic hierarchical category in the dataset. In Table 10, we compare the performance of our method with that of these two approaches. Regarding the CHiLS experiment, as their paper indicates that it is tested only on the 20 classes of CIFAR-20 and 101 classes of Food-101, to maximize the representation of its effect, we follow its setup and do not test it on the remaining hierarchical categories. The results show an improvement on the generic image dataset, but a decline on the fine-grained food dataset, consistent with our previous analysis. In the Hierarchy-CLIP experiment, it also exhibits commendable performance, showing improvements across various hierarchical levels. This is largely attributed to its use of both superclasses and subclasses of labels for expansion. Our method not only enhances classification performance at each level but also can be conveniently transferred to downstream datasets.
F.3 Params number and FLOPs
Table 11 shows the computational complexity of HGCLIP in comparison with previous works. Although HGCLIP leverages graph encoders, its overall FLOPs are only 0.1% higher than MaPLe, and it offers better convergence and competitive training/inference speed.
Type Method Params
Params
%CLIP
GFLOP
Train time
FPS Acc. %
Prompt Learning CoOp Zhou et al. (2022b) 2k 0.002 163.4 13.52 1345 80.29
CoCoOp Zhou et al. (2022a) 36k 0.03 163.4 42.18 15.3 79.17
MaPLe Khattak et al. (2023a) 3.55M 2.85 192.7 22.34 1357 88.24
PromptSRC Khattak et al. (2023b) 32k 0.02 181.1 15.43 1379 85.21
[…]
Tip-Adapter Zhang et al. (2022) 0.4k 3e-6 203.0 7.16 1466 83.00
[…]
FGN Chang et al. (2021) 29.3M - 120.17 78.57 38.04 85.74
CHRF Liu et al. (2022) 31.6M - 118.42 89.15 33.50 86.79
TFGIC Xu et al. (2023) 31.8M - 119.56 88.06 36.36 87.19
[…]
HGCLIP 4.25M 3.41 193.6 20.17 1349 89.21
Table 11: Computing cost comparison on CIFAR-100 dataset. $\dagger$ denotes the GNN parameters of the two modalities are shared. FPS denotes Frames Per Second (with a batch size of 100).